OpenStudy (anonymous):

Find and simplify the difference quotient f(x + h) − f(x)/h for the following function: f(x) = x2 − 9x + 8

OpenStudy (anonymous):

which part are your having trouble with?

OpenStudy (anonymous):

all of it, i dont understand how to do it

OpenStudy (anonymous):

f(x+h) is f(x) but with (x+h) plugged in as every x so f(x+h) = (x+h)^2 - 9(x+h) + 8 you'll want to expand that and find [f(x+h) - f(x)]/h

OpenStudy (anonymous):

so itd be x^2+h^2-9x-9h+8 ?

OpenStudy (anonymous):

????

OpenStudy (anonymous):

(x+h)^2 = x^2 + 2xh + h^2 but the rest is correct

OpenStudy (anonymous):

okay but could you explain how you get 2xh?

OpenStudy (anonymous):

when multiplying two binomials, you use FOIL: [steps 1,2,3,4]